### v3

# Weak-Strong Uniqueness and Extreme Wall Events at High Reynolds Number

Abstract
Singular or weak solutions of the incompressible Euler equations have been hypothesized to account for anomalous dissipation at very high Reynolds numbers and, in particular, to explain the d’Alembert paradox of non-vanishing drag. A possible objection to this explanation is the mathematical property called “weak-strong uniqueness”, which requires that any admissable weak solution of the Euler equations must coincide with the smooth Euler solution for the same initial data. As an application of the Josephson-Anderson relation, we sketch a proof of conditional weak-strong uniqueness for the potential Euler solution of d’Alembert within the class of strong inviscid limits. We suggest that the mild conditions required for weak-strong uniqueness are, in fact, physically violated by violent eruption of very thin boundary layers. We discuss observational signatures of these extreme events and explain how the small length-scales involved could threaten the validity of a hydrodynamic description.
pacs:
?????I Introduction
Various empirical observations suggest that drag and dissipation are non-vanishing for turbulent flows of incompressible fluids at low Mach number, even in the limit of infinite Reynolds number. Evidence provided by laboratory experiments necessarily involves flow interactions with solid surfaces, such as asymptotic non-zero values of drag coefficients for flow past bodies: Frisch (1995), §5.2. It was conjectured by Onsager Onsager (1949); Eyink (2024a) that such turbulent “anomalous dissipation” can be explained by singular or weak solutions of ideal Euler equations and recent works have developed this theory for flows interacting with solid walls. In particular, a solution of the famous d’Alembert paradox d’Alembert (1749, 1768) can be formulated within this approach, according to which the drag on the body in the infinite Reynolds number limit is produced by some weak Euler solution with vorticity in the wake, supplanting the potential flow of d’Alembert that exerts no drag Hoffman and Johnson (2010); Quan and Eyink (2024).
A challenge for this proposed solution is the notion of “weak-strong uniqueness” in the mathematical theory of partial differential equations Wiedemann (2018). Since this concept seems to be not widely known to fluid physicists, we shall summarize it here succinctly and explain the difficulty it poses for Onsager’s theory. We shall also sketch a simple proof for the case of the d’Alembert flow Quan and Eyink (2025a), which exploits the Josephson-Anderson relation recently derived for such external flows Eyink (2021, 2024b). The crucial point of our work is that weak-strong uniqueness holds only conditionally for fluids interacting with solid boundaries. If these sufficient conditions do not hold, then the challenge to Onsager’s theory from weak-strong uniqueness is eliminated. The sufficient conditions appear rather mild, so that, conversely, any violation of weak-strong uniqueness requires quite extreme events. In fact, we shall argue that these events are so singular that they could threaten the very validity of a continuum hydrodynamic description. It is the main purpose of this paper to delineate the striking signatures of this breakdown in order to facilitate further investigations by laboratory experiment and by numerical simulation. Our discussion may be useful also for mathematicians as a short summary of relevant empirical observations.
II What is Weak-Strong Uniqueness?
An informal statement of weak-strong uniqueness is as follows: If a smooth Euler solution exists on spacetime domain with initial data , then any generalized Euler solution on the same domain with the same initial data must coincide with that smooth solution, so long as the generalized solution satisfies some modest “admissibility” requirement. A common example of an admissibility requirement is the global energy bound
which is an expected consequence of energy dissipation. In the spatial domains or without boundaries for , any locally dissipative weak Euler solution (in particular any weak solution obtained as the strong inviscid limit of Navier-Stokes solutions) enjoys the weak-strong uniqueness property. Such strong limits exist under certain reasonable conditions Drivas and Nguyen (2019) but even more generalized notions of “Euler solution” are guaranteed to exist in the inviscid limit (at least along a subsequence of viscosities) and also enjoy weak-strong uniqueness. These include dissipative Euler solutions “in the sense of Lions” Lions (1996); Bardos and Titi (2013) and measure-valued Euler solutions Brenier et al. (2011). See Wiedemann (2018) for a very lucid mathematical review of these and related results. A common view expressed in the mathematics literature is that weak-strong uniqueness is a necessary requirement for any “reasonable” notion of generalized Euler solution. The underlying presumption is that a smooth Euler solution, whenever it exists, must be the “physical” solution in the inviscid limit.
One important implication of weak-strong uniqueness is on existence of blow-up or finite-time singularities for smooth Euler solutions. Indeed, weak-strong uniqueness for the inviscid limit solution implies that viscous energy dissipation must necessarily vanish over any finite time interval for which the smooth Euler solution exists Eyink (2008); Bardos and Titi (2013). Conversely, appearance of anomalous energy dissipation over a finite time interval for a flow with smooth initial data requires that the incompressible Euler solution with that initial data cannot remain smooth over the interval. A typical example of such initial data is the Taylor-Green vortex in periodic domain Taylor and Green (1937), although existing evidence for anomalous dissipation in this flow Fehn et al. (2022) seems less compelling to us than corresponding evidence for wall-bounded flows. In any case, a finite-time Euler singularity can be rigorously ruled out for certain smooth initial data, e.g. for any smooth data in a 2-dimensional spatial domain and for the stationary potential data in the 3-dimensional d’Alembert flow. Thus, finite-time blow-up of smooth Euler solutions cannot be the general route to turbulence and anomalous dissipation in incompressible fluid flows.
III The New D’Alembert Paradox
Perhaps the most famous example of a smooth Euler solution which exists globally in time in a three-dimensional space domain is the solution of d’Alembert for potential flow past a smooth body. The context is a rigid body occupying a closed set with a smooth boundary immersed in a fluid filling an infinite-volume domain with velocity at infinity. Alternatively, by Galilean invariance, one may consider a rigid body moving with translational velocity through a fluid at rest at infinity, but we find it more convenient to discuss the problem in the rest frame of the body. There is a unique potential flow solution of the incompressible Euler equation
where the the potential is defined at each time instant as the solution of the Neumann boundary value problem
unique up to a spatial constant, and with pressure likewise determined up to a spatial constant from the Bernoulli equation
If the velocity is smooth as function of time, then this unique potential Euler solution is smooth on the space-time domain
The force exerted on the body by the potential Euler flow past it is given instantaneously by the integral
where is the unit normal on the body surface pointing into the fluid. For simplicity, we have taken fluid density here and throughout. The famous result of d’Alembert d’Alembert (1749, 1768) for the stationary potential flow with constant fluid velocity is that the force on the body vanishes identically, This result can be generalized to time-dependent potential flow as long as is bounded in time and, thus, the fluid impulse Lighthill (1986)
is also bounded. Since it then follows easily that long-time average force must vanish:
In addition, the power dissipated by fluid drag force, can be written likewise as a total time-derivative by noting that in terms of the “added mass tensor” of the rigid body Lighthill (1986). In that case, and one similarly obtains
This vanishing seems to contradict common experience that drag forces are non-zero and it led to the famous “d’Alembert paradox.” The accepted resolution Stewartson (1981) of this original paradox is that proposed by Saint-Venant and Navier, which is that molecular fluids have a viscosity and are governed by the Navier-Stokes equation
which predicts a non-vanishing time-average drag dissipation for any finite value of the Reynolds number Here is diameter of the body
However, a new d’Alembert paradox arises if one considers the limit Stewartson (1981). There is substantial empirical evidence that drag remains non-vanishing in this limit, for example, the apparent non-zero asymptotic values of drag coefficients for solid bodies of various shapes. Nevertheless, any such limits must be described by generalized Euler solutions Lions (1996); DiPerna and Majda (1987) or, under reasonable assumptions, by weak Euler solutions as conjectured by Onsager Drivas and Nguyen (2019) and a suitable generalization of the admissability condition (II.1) furthermore must hold. Weak-strong uniqueness would require that the generalized Euler solution obtained in the limit must coincide with the smooth Euler solution of d’Alembert, suggesting that time-average drag must vanish in the limit To argue for this conclusion, one may consider the situation where the fluid starts at rest with and then smoothly accelerates to a velocity bounded in time. Taking so large that , then weak-strong uniqueness implies Quan and Eyink (2025a) that
Note, however, that tabulated drag coefficients are not so well-characterized experimentally, e.g. the body may have accelerated through fluid with small turbulent fluctuations and not initially at rest. Thus, it is not completely clear that weak-strong uniqueness contradicts the observations of non-vanishing drag.
Beyond simple drag coefficients, however, more carefully controlled laboratory and numerical experiments have been performed to measure force histories and other characteristics for the case of accelerated bodies. This is particularly true for the classical problem of the impulsively accelerated cylinder initiated in the 1925 work of Prandtl Prandtl (1925) and since extensively studied experimentally Prandtl and Tietjens (1931); Schwabe (1935); Honji and Taneda (1969); Taneda (1977); Bouard and Coutanceau (1980); Loc and Bouard (1985); Nagata et al. (1985a, b) and computationally Payne (1958); Collins and Dennis (1973); Chang and Chern (1991); Koumoutsakos and Leonard (1995); Mittal et al. (2008); Chatzimanolakis et al. (2022). This case is interesting because it corresponds in the body frame to the Navier-Stokes solution with initial data given by d’Alembert’s stationary potential flow, . Notice for this initial data that corresponds to a singular vortex sheet at the body surface. A delicate aspect of such impulsively accelerated flows is that the very earliest stages cannot be described hydrodynamically. Kinetic theory shows that slightly modified initial conditions and small slip at the boundary are necessary to match the solutions of Boltzmann equation and Navier-Stokes equation. These complications make the comparison of the Prandtl problem with hydrodynamic theory a bit more involved, as we discuss carefully in Section V. It is easier to analyze theoretically the case of gradually accelerated bodies, a case which has been studied also quite extensively. We mention here just a few papers on experiment Taneda and Honji (1971); Pullin and Perry (1980); Grift et al. (2019); Reijtenbagh et al. (2023) and simulation Koumoutsakos and Shiels (1996); Homann et al. (2013); Xu and Nitsche (2015); Sader et al. (2024), out of a huge literature. None of these various studies show any clear evidence of the high-Reynolds number flow converging to d’Alembert’s potential flow solution. Although further careful studies at much higher Reynolds numbers are required to provide more convincing data, we shall explain next a physical mechanism that can violate weak-strong uniqueness and thus permit a singular Euler solution in the inviscid limit distinct from d’Alembert’s smooth potential solution.
IV Weak-Strong Uniqueness Can Fail With Solid Walls
The traditional view in the mathematics community has been that weak-strong uniqueness is a necessary requirement for any “reasonable” concept of generalized solution, but recent mathematical works on flows with solid boundaries have called into question this presumption. A sigificant step in this direction is the result on non-uniqueness of weak Euler solutions in an annular domain for piecewise-smooth initial data with an interior vortex sheet, established by convex integration methods Bardos et al. (2014). The initial data in this example is a stationary strong Euler solution, which co-exists with infinitely many admissible weak solutions that exhibit spreading mixing layers evolving from the initial vortex sheet. This is not a perfect counterexample to weak-strong uniqueness, however, because the stationary Euler solution has a jump discontinuity in its velocity field. A much cleaner example has been established recently, also by exploiting convex integration methods developed for vortex-sheet initial data Székelyhidi Jr (2011). In the case of a plane-parallel channel with initial data given by constant plug flow, , it has been proved that the obvious stationary solution of Euler (, plug flow) co-exists with infinitely-many admissible weak Euler solutions in which the surface vortex sheet separates and mixes in the interior Vasseur and Yang (2023). Since plug flow is (in fact, analytic), this is a clear mathematical counterexample showing how weak-strong uniqueness can be violated by separation at the boundary. Such weak Euler solutions with boundary separation may not be physically realizable in channel flow with smooth, plane-parallel walls, as we discuss in Section VII.1 below, but we argue that weak-strong uniqueness can be violated by this mechanism in physical examples such as the d’Alembert potential flow.
Although clear counterexamples were found only recently, it has been known for awhile that standard proofs of weak-strong uniqueness break down in flows with solid walls unless additional conditions on the solutions beyond (II.1) are imposed. For example, Bardos and Titi (2013) have shown that weak-* limits of Navier-Stokes solutions in a bounded domain with stick boundary conditions are dissipative Euler solutions in the sense of Lions on and thus satisfy weak-strong uniqueness, but only under the following condition: the wall shear stress or skin friction on the boundary must satisfy
with convergence in the sense of distributions on the domain The same result was established in Bardos and Titi (2013) for Navier-slip boundary conditions as well, as long as the normalized slip length as It was also proved in Bardos and Titi (2013) that (IV.1) is implied by the condition of vanishing dissipation in a neighborhood of the boundary
where the boundary layer of thickness is defined by
This is the famous condition shown by Kato Kato (1984) to be equivalent to strong convergence of the Navier-Stokes solution to a strong Euler solution, when the initial data also converge strongly and as long as the strong Euler solution exists. Since such strong limits of Navier-Stokes solutions are admissible weak Euler solutions, Kato’s theorem can be considered one of the earliest results of weak-strong uniqueness type. A more recent work of Kelliher Kelliher (2017) has comprehensively discussed these various conditions for inviscid limits of closed flows in spatially bounded domains and shown, in particular, that the condition (IV.1) can be weakened to
where is an assumed strong Euler solution. Another class of conditions sufficient for weak-strong uniqueness involve some continuity of velocity for the admissible weak Euler solution near the boundary. This was first proved by Bardos et al. (2014) who required that the velocity field be Hölder continuous in some neighborhood with and then proved by Wiedemann (2018) assuming only continuity in . Recently, we have shown Quan and Eyink (2025a) that even less stringent conditions on the weak Euler solution are sufficient for weak-strong uniqueness within the general class of admissible weak Euler solutions, namely:
for some or near-wall boundedness, and further
Note that where the “essential supremum” is the least upper bound of values such that the set has positive measure. The latter condition (IV.5) can be interpreted as uniform continuity at the boundary of the normal velocity component, since We refer to (IV.4)-(IV.5) as the “Drivas-Nguyen conditions” since they were first employed in Drivas and Nguyen (2018) to study anomalous energy dissipation in wall-bounded flow.
Although these conditional weak-strong uniqueness results have previously been interpreted as identifying additional “admissibility criteria” for weak Euler solutions Wiedemann (2018), we argue below that weak-strong uniqueness probably fails for the physical weak solutions and thus all of these apparently modest conditions must be violated. A main example that motivates our claim is the d’Alembert potential Euler flow, which we next discuss.
V Proof Sketch for D’Alembert Flow
To our knowledge, no prior theorem on conditional weak-strong uniqueness has been applicable to the d’Alembert potential Euler solution, until we recently established such a result Quan and Eyink (2025a). The theorems of Bardos and Titi (2013) on inviscid limits apply to flow domains of unbounded extent but assume that the smooth Euler solution has finite energy, which is untrue of the d’Alembert flow with a non-vanishing velocity at infinity. Our result in Quan and Eyink (2025a) is likewise proved for inviscid limits where the initial data for the Navier-Stokes solution converge strongly in to the potential flow with velocity at infinity. As we discuss below, this assumption is true both for the case of impulsive acceleration and for gradual acceleration from rest. We assume finite kinetic energy (spatial norm) only for the rotational fluid velocity which is realistic because this field is dipolar at infinity Eyink (2021). The main technical tool that we employ is the Josephson-Anderson relation for drag on the body Eyink (2021), which appears as a source term in the balance equation for kinetic energy of rotational fluid motions in the wake:
We thus use a version of a standard “relative energy” argument to prove weak-strong uniqueness Wiedemann (2018). In this context, our sufficient condition for weak-strong uniqueness is that for all times, which is nearly the same condition (IV.3) invoked by Kelliher for closed flows in bounded domains.
Because the essence of the proof in Quan and Eyink (2025a) is quite simple, it is instructive to sketch the main details here. The starting point of the proof is the global balance for the rotational kinetic energy in the closed domain so that for a.e.
Here is the anomalous energy dissipation and the last two terms are the inviscid limit of the Josephson-Anderson relation for power dissipated by drag, which appears as a source term in the rotational energy balance. Much of the technical work of the proof involves the justification of the energy balance equation (V.2), by smearing a corresponding local balance equation with a test function which is a smoothed version of the characteristic function where is the ball of radius centered at the origin, restricted to and then taking the limit
Assuming that , then since
where the last line follows by Cauchy-Schwartz inequality. Finally, by (V.4) and the Gronwall inequality, one obtains for a.e.
Thus, when so then it follows that
This argument shows that any strong- inviscid limit with initial data must in fact coincide with for a.e. times . The applicability of this result to a smoothly accelerated body with is straightforward, assuming the validity of the incompressible Navier-Stokes model, since The relevance to the impulsively accelerated body requires more detailed justification because of the singular nature of that problem. It is worthwhile to discuss here since it raises some physics issues which will be important later.
Our derivation of the inviscid Josephson-Anderson relation Quan and Eyink (2024) and rigorous proof of weak-strong uniqueness for the d’Alembert potential Euler solution Quan and Eyink (2025a) both assume existence of a strong Navier-Stokes solution on the domain for any viscosity However, this assumption cannot be true up to time for the initial data because and the first compatibility condition for a continuous solution is thus violated Temam (2006). However, the Navier-Stokes solution for this initial data can be expected to be smooth on the time-interval for any time Temam (2006). Indeed, for the Prandtl problem of an impulsively accelerated cylinder, a formal Navier-Stokes solution has been obtained for at short dimensionless times with fixed but arbitrary Bar-Lev and Yang (1975). The non-dimensionalization here employs outer units, with lengths normalized by the cylinder radius velocities normalized by the flow velicity at infinity, and times normalized by The formal solution is obtained by a third-order matched asymptotic expansion in which the outer potential-flow solution to leading order is the d’Alembert solution and the inner solution describes a viscous boundary layer of thickness at the cylinder suface. Although this formal solution has never been rigorously derived, to our knowledge, it agrees well with high-resolution numerical simulations of the Prandtl problem at early times Koumoutsakos and Leonard (1995); Chatzimanolakis et al. (2022). Since the boundary layer of thickness corresponds to with vanishingly small kinetic energy as we may consider instead the Navier-Stokes solution on the time interval with for some fixed and it remains true that strongly in
The above mathematical explanation nevertheless suffers from a physical defect, because the formal asymptotic solution has a skin friction diverging as at due to the initial singular vortex sheet at the body surface Bar-Lev and Yang (1975). This solution for impulsive acceleration of the body is obviously experimentally unrealizable. If the acceleration of the body really occurs on a time scale of order the mean-free collision time of the molecular fluid or on an even shorter time scale, then the problem can no longer be described accurately by Navier-Stokes equation with stick boundary conditions. A more detailed analysis by methods of kinetic theory Reddy (1967) suggests that one may still obtain a uniformly accurate solution by Navier-Stokes equations, if one replaces the stick boundary conditions with Navier-slip conditions for a slip length of order the molecular mean-free-path length and if the initial data are modified by a “kinetic boundary layer” or Knudsen layer with thickness also of order . The latter layer removes the divergence in the skin friction at , as physically required. Note that this boundary layer in dimensionless outer units has thickness with the Mach number, and the kinetic description modifies the previous asymptotic Navier-Stokes solution at very short times We used here the standard estimate from kinetic theory for kinematic viscosity with the mean-free-path length and with the sound speed. The drag coefficient no longer diverges at but instead assumes a large value Reddy (1967). The proof of weak-strong uniqueness for the d’Alembert potential flow given in Quan and Eyink (2025a) works with Navier-slip b.c. as long as the slip length in dimensionless outer units vanishes as See Appendix A. Thus, our mathematical results still apply in this kinetic reformulation.
Some doubts remain, however, whether our analysis applies to a physical realization of impulsive acceleration in a laboratory experiment. The kinetic description of a gas by the Boltzmann equation in the low-density limit is now believed to be incomplete, missing stochastic effects of molecular fluctuations Spohn (1981); Bouchet (2020); Sawant et al. (2021). Such fluctuation effects do not necessarily vitiate Onsager’s theoretical description of high Reynolds number flows by weak Euler solutions Eyink and Peng (2024). However, to our knowledge, weak-strong uniqueness results have not yet been proved for inviscid limits of models incorporating fluctuations. Thus, it seems advisable to focus theoretical treatment and future empirical investigation on the more regular (and commonplace) problem of gradually and smoothly accelerated bodies. Rather surprisingly, however, we shall find that even the problem of smooth acceleration may encounter such difficulties, because breakdown of weak-strong uniqueness requires such extreme events that the validity of a macroscopic hydrodynamic description again is threatened. See Section VII.3.
VI Two Alternative Scenarios
Assuming validity of a hydrodynamic description by Navier-Stokes, our mathematical results lead to two distinct alternatives, both of which might be argued to be consistent with empirical observations of turbulent drag on solid bodies. We discuss these two alternatives in turn.
VI.1 First Alternative:
Under this condition, weak-strong uniqueness holds. As a consequence, for initial data strongly converging to pure potential, drag on a body vanishes over any fixed time-interval as see (III.4). If the drag vanishes sufficiently slowly, however, then it might be very difficult to distinguish empirically from drag strictly non-vanishing. The phenomenon of very slowly vanishing drag and dissipation has been termed a “weak dissipative anomaly” Bedrossian et al. (2019). Such weak anomalies seem to occur physically in a number of wall-bounded flows, such as circular pipe flows with perfectly smooth walls (see Eyink (2024a) for a summary). Convergence to a smooth potential Euler solution has been suggested to occur also in some bluff body flows, e.g. rapidly rotating cylinders where separation is argued to be suppressed Glauert (1957), although recent high Reynolds number simulations do not seem to support that conclusion Aljure et al. (2015). Strong anomalous dissipation and drag may furthermore occur experimentally if the initial data is not pure potential in the high-Reynolds number limit. Although the rotational flow from a thin viscous boundary layer vanishes in the energy norm as there may in addition be small but non-vanishing vorticity in the incoming flow. This is similar to what occurs in so-called “bypass transition”, where small levels of turbulent fluctuations in the background flow can trigger laminar-to-turbulent transition. In order to discuss both of the above possibilities more quantitatively, we can apply the analysis of the previous section.
The first possibility of “weak anomaly” is best discussed by generalizing this analysis to finite Reynolds number or equivalently positive viscosity . From the global energy balance for the rotational motions, one obtains
by using and Cauchy-Schwartz inequality. Then by Gronwall inequality, we obtain
Even though the prefactor of the exponential vanishes in the limit with fixed, the upper bound does not generally vanish for fixed as Consider the case where the initial rotational flow arises entirely from a viscous boundary layer of vanishingly small thickness for with In that case, as and, by our main assumption in this section, also However, when (as for a Prandtl boundary layer with ), then the Reynolds number of the boundary layer itself will be very large for and thus prey to various instabilities. The successive instabilities of Prandtl layers and subsequent thinner sublayers have been studied both by fluid mechanicians Cassel and Conlisk (2014) and by mathematicians Grenier and Nguyen (2018, 2019). Such instabilities provide a mechanism for generation and growth of rotational flow, so that
Even in very quiet flow without tiny external perturbations, intrinsic thermal noise might trigger instability and flow separation McMullen and Gallis (2024).
It is important in this context to distinguish between two different notions of stability. On the one hand, there is the concept of stability in mechanics or dynamical systems, according to which a solution is “stable” if small perturbations do not grow or even decay in magnitude. In the opposite case where infinitesimal perturbations grow in magnitude, the solution is called dynamically unstable. On the other hand, another notion of stability of solutions in applied mathematics is “well-posedness”, which holds when existence, uniqueness and continuity in the data (initial conditions, equations of motion, etc.) are all guaranteed. This property is sometimes called “Hadamard stability” after the mathematician who first codified the concept Hadamard (1902). In the opposite case, the solution is called ill-posed or Hadamard unstable. What must be emphasized is that Hadamard stability is generally a much weaker requirement than dynamical stability and it is perfectly consistent with exponential growth of small perturbations. For example, even smooth chaotic dynamical systems in which every solution exhibits exponential sensitivity to initial data are well-posed in the sense of Hadamard.
Weak-strong uniqueness is precisely a statement of well-posedness of the classical smooth Euler solution, i.e. its Hadamard stability, even within a much larger class of “admissable, generalized Euler solutions”. This point is made very clearly by the basic inequality (V.5). This result implies that any “viscosity solution” of Euler, obtained by a strong inviscid limit, must coincide with the smooth potential solution of d’Alembert, if or In addition to this uniqueness statement, one can infer also from (V.5) continuity in the initial data: any “viscosity solution” can be made to agree arbitrarily closely with over any finite time interval if or As we have just discussed, however, this “Hadamard stability” of the d’Alembert solution is perfectly compatible with its dynamical instability. This fact is directly relevant to the situation where the initial data for Navier-Stokes do not converge strongly in to because turbulent fluctuations with some small but non-vanishing energy are superimposed on the potential flow. At long times the viscosity solution with such initial data can depart far from the d’Alembert solution and can exhibit a strong dissipative anomaly.
These conclusions have implications for some interesting proposals of Hoffman & Johnson on the solution of the d’Alembert paradox Hoffman (2006); Hoffman and Johnson (2007, 2010). In agreement with the earlier ideas of Onsager, they explain non-vanishing or anomalous drag via “turbulent Euler solutions”, which they have attempted to calculate from numerical finite-element schemes. They summarize their proposed solution as follows:
“We have presented a resolution of d’Alembert’s Paradox based on analytical and computational evidence that a potential solution with zero drag is illposed as a solution of the Euler equations, and under perturbations develops into a wellposed turbulent solution with substantial drag in accordance with observations” Hoffman and Johnson (2010).
Recognizing that such weak Euler solutions may be non-unique even for exactly specified initial data the authors suggest an interesting concept of “output well-posedness” according to which some space-time averaged outputs (such as mean drag) may be unique and continuous in the inital data, even though individual Euler solutions are not. These proposals are generally consistent with our rigorous mathematical analysis, with just one important exception: the d’Alembert potential flow can be dynamically unstable but it cannot be, as claimed by Hoffman and Johnson (2010), “ill-posed as a solution of the Euler equations,” at least not when weak-strong uniqueness holds. As we discuss in the following section, and at length in Section VII, ill-posedness of the d’Alembert potential flow in fact requires very extreme wall events with quite striking experimental signatures.
VI.2 Second Alternative:
When the condition is not valid identically in spacetime (in the sense of distributions), then weak-strong uniqueness may not hold for the d’Alembert potential flow within the class of strong inviscid limits. In particular, the potential Euler solution may co-exist with other weak Euler solutions obtained via inviscid limits, with identically the same initial data, but with non-vanishing vorticity and drag! This scenario recalls the result for plane-parallel channel geometry of Vasseur and Yang (2023) that plug flow coexists with infinitely-many, admissible weak Euler solutions exhibiting separation. As we discuss in the next Section VII, available evidence suggests that the conditions for weak-strong uniqueness in fact hold in channel flows. On the other hand, we shall argue that breakdown of weak-strong uniqueness is a very plausible possibility for flows around solid bodies, permitting coexistence of the smooth d’Alembert flow and dissipative Euler flows with separation and drag. We discuss next the extreme flow events that are required for such breakdown to occur.
VII Extreme Wall Events
We have already discussed in Section IV three rather modest-looking conditions that suffice for weak-strong uniqueness. Within the class of weak-* limits of Navier-Stokes solutions with bounded energy in general domains, Bardos and Titi (2013) have shown that the condition (IV.1) of vanishing skin-friction suffices for weak-strong uniqueness of any smooth Euler solution. Furthermore, Bardos and Titi (2013) have shown that the condition (IV.3) of vanishing energy dissipation in a “Kato layer” implies the previous condition (IV.1). In fact, the third set of conditions (IV.4),(IV.5) on continuity of normal velocity at the wall implies also the vanishing skin-friction condition (IV.1) for strong inviscid limits of Navier-Stokes solutions. The proof in Quan and Eyink (2025b) is based on the analogy between skin friction and viscous energy dissipation, on the one hand, and energy cascade through scales and momentum cascade through space, on the other hand Jiménez (2012). Thus, non-vanishing skin friction can be understood as a “strong momentum anomaly” and the continuity conditions (IV.4),(IV.5) are analogous to the Onsager condition on velocity Hölder exponent which forbids an energy to small scales. The analogous idea of the proof in Quan and Eyink (2025b) is that non-vanishing velocity toward the wall, at any positive distance, is required to carry momentum to the viscous sink at the wall in the infinite Reynolds number limit. Thus, a strong momentum anomaly is possible only if the normal velocity is is not uniformly continuous at the wall.
Somewhat surprisingly, the conditions for a momentum anomaly and breakdown of weak-strong uniqueness are much more severe than those needed for energy cascade, requiring essentially a shock-like “discontinuity” at the wall, or an Hölder singularity! Interestingly, this possibility seems to have been anticipated by G. I. Taylor as early as 1915, who wrote the following in discussing interactions of atmospheric turbulence with the solid ground:
“…a very large amount of momentum is communicated by means of eddies from the atmosphere to the ground. This momentum must ultimately pass from the eddies to the ground by means of the almost infinitesimal viscosity of the air. The actual value of the viscosity of the air does not affect the rate at which momentum is communicated to the ground, although it is the agent by means of which the transference is effected.
The finite loss of momentum at the walls due to an infinitesimal viscosity may be compared with the finite loss of energy due to an infinitesimal viscosity at a surface of discontinuity in a gas.” Taylor (1915)
The Kato condition (IV.3) likewise involves an extreme situation of non-vanishing viscous dissipation in a shock-like layer at a wall. Kato-type boundary layers of thickness it is worth noting, were proposed already in 1923 by Burgers as a mechanism to produce anomalous drag Burgers (1923)111Burgers’ discussion was longer than Taylor’s Taylor (1915) but his arguments, as he admitted himself, were “in many respects vague”. He proceeded by deriving integral mean balances for momentum and energy and by investigating the parameter values to maximize drag. He correctly concluded that a boundary layer with thickness would produce skin friction non-vanishing as Unfortunately, this argument leads to non-vanishing drag coefficient for smooth walls, whereas Burgers had noted himself that only “for rough-walled tubes is approximately independent” of .. The corresponding rigorous conditions for weak-strong uniqueness, although originally derived for internal flows in bounded domains, apply also to external flows around bodies: see Section V.
The severe events required to violate weak-strong uniqueness possess very striking signatures that should be observable empirically, if they exist, both in numerical simulations and in laboratory experiments. There is already very active current investigation of extreme events in wall-bounded turbulence. However, the events required to violate weak-strong uniqueness are far more violent than those reported in the prior literature, with one prominent exception Nguyen van yen et al. (2018). This relative lack of evidence may be due to the dominant focus of such studies on the “canonical wall-bounded flows”: plane-parallel channel flow Lenaers et al. (2012); Srinivas et al. (2024), circular pipe flow Guerrero et al. (2020); Fei et al. (2024), and the flat-plate boundary layer Pan and Kwon (2018). Although these flows are often considered to capture the essence of turbulent-solid surface interactions in the simplest geometry, in fact these flows exhibit a number of very atypical features, as a direct consequence of their simplicity. For example, none of these canonical flows show any evidence for a strong energy dissipation anomaly, in contrast to common wake flows past solid bodies or internal flows with hydraulically rough walls Eyink (2024a). We summarize below current observations on extreme events, first for channels as representative of “canonical flows” and then for the single flow that, to our knowledge, provides positive evidence.
VII.1 Channel Flow
The relevant quantities for weak-strong uniqueness are skin friction, wall-normal velocity and near-wall viscous dissipation, which we discuss in turn.
VII.1.1 Skin Friction
The condition may seem improbable, given the very common observation that time-average skin friction vanishes in the high Reynolds limit, as This observation even holds when drag is non-vanishing, as for bluff bodies Achenbach (1972) and rough pipes Busse et al. (2017), because the asymptotic net drag in those instances is apparently supplied entirely by pressure forces (form drag). However, in order to violate weak-strong uniqueness it is enough that not vanish over some finite region of space-time and it need not be true that any non-zero fraction of the asymptotic drag arises from skin friction. Thus, rather than the mean value, it is more appropriate to consider the entire distribution of values over the wall at long times.
The probability distribution of wall shear stress has been a focus for previous numerical studies Lenaers et al. (2012); Srinivas et al. (2024), in particular the streamwise component which contributes to drag. Note that it is also non-vanishing of the streamwise component which is relevant for possible physical violation of weak-strong uniqueness for plug-flow in a channel, with a spacetime constant velocity in the streamwise or -direction. In fact, since all of the theorems about infinite- limit cited in this paper require non-dimensionalization of flow variables in outer units, what is relevant is scaled by Since the results in Lenaers et al. (2012); Srinivas et al. (2024) are instead scaled by the friction-velocity squared, we have scanned the data in Figure 2 of Lenaers et al. (2012) and rescaled the results by the factor The parameterization of the Prandtl-Kármán drag law from equation (12) of Zanoun et al. (2021) was used to determine from and hence the ratio The results for the probability density function of in outer units are plotted in Figure 1, at three values of the Reynolds number
The trends with increasing Reynolds number are clear. The mean values are decreasing very slowly, consistent with the logarithmic decay predicted by the Prandtl-Kármán law. Simultaneously, the widths of the probability distributions are shrinking slowly also with increasing This decrease in widths is a consequence of our scaling in outer units, whereas prior observations of the probability distribution of skin friction in inner units revealed increasing variances and increasing far tails as Reynolds numbers rose Lenaers et al. (2012). Although the convergence suggested by our Figure 1 is quite slow, the results are consistent with as According to this empirical data, weak-strong uniqueness is likely to hold for plug flow through the channel, within the class of inviscid limits.
VII.1.2 Wall-Normal Velocity
The Drivas-Nguyen conditions for weak-strong uniqueness involve, in particular, extreme values of the wall-normal velocity (essential suprema) which are non-vanishing approaching the wall. Extremes of the wall-normal velocity have previously been studied in channel flow, for example, see Lenaers et al. (2012), Figure 5, for probability distributions of the wall-normal velocity. However, those distributions are for wall-normal velocities scaled by their standard deviation and involve positions only in the viscous sublayer and buffer layer. Instead, the Drivas-Nguyen condition involves -extremes in the inertial layer where direct viscous effects are entirely absent. In dimensional terms, the Drivas-Nguyen condition on wall-normal velocity is that
for some Thus, the maximum of for distances must remain for some as through the inertial layer. The maximization can presumably be restricted to the inertial interval (in inner units, ), since the largest values should occur there. There have been some previous studies of extreme values in the logarithmic layer of turbulent channel flow, but for viscous dissipation rather than velocities Hack and Schmidt (2021). There is some indication from Figure 6 of Lenaers et al. (2012) that extremes of wall-normal velocity become highly improbable at distances if an “extreme value” is defined as there to be a value greater than 10 standard deviations from the mean. However, note that the condition (VII.1) requires instead velocities only some small fraction of but arbitrarily close to the wall. We are aware of no direct evidence for this condition in channel flow, but it is not implausible.
VII.1.3 Dissipation in a Viscous Near-Wall Layer
The Kato condition (IV.3) for channel flow is just the requirement that the net energy dissipation integrated over the viscous sublayer and buffer layer should vanish relative to as Although there have been a great many studies of viscous dissipation in turbulent channel flows, we are not aware of any study of this particular issue. The prediction of the Prandtl-Kármán theory is that dissipation in the buffer layer and viscous sublayer should scale as smaller than the total dissipation in the log-layer by a factor of about Both, however, tend to zero relative to as These considerations suggest that the Kato condition should be satisfied and indeed convergence to plug flow has been argued to occur in smooth-walled, plane-parallel channel flows Cantwell (2019).
If so, then the counterexamples of Vasseur and Yang (2023) to weak-strong uniqueness in channel flows are of mathematical relevance only and would not appear in the inviscid limit. Note, however, that all of the observations that we have discussed are for statistically stationary, fully-developed turbulent channel flow. Weak-strong uniqueness involves instead the initial-value problem starting with plug flow, which may be achieved, for example, by impulsive acceleration of the channel walls to velocity . We are not aware of any relevant empirical studies in this context.
VII.1.4 Vortex-Induced Separation
Although most available observations on channel flow do not involve the initial transient regime and are thus not directly relevant to the issue of weak-strong uniqueness, it is nevertheless of some interest to understand what fluid mechanical events lead to the most extreme near-wall events in fully-developed channel flow. The events which produced large wall-normal velocities (both positive and negative) were visualized in Figure 9 of Lenaers et al. (2012) and as a composite image obtained by conditional averaging in figure 10 of of Lenaers et al. (2012). These results, confirmed by subsequent studies, show that extreme wall-normal velocity events in the viscous sublayer and buffer layer of turbulent channel flow with smooth walls have the general characteristics of vortex-induced separation Doligalski et al. (1994). These extreme events are triggered by strong quasi-streamwise vortices which approach near the wall and induce very strong motions both toward and away from the wall. Extreme values of wall-stress Sheng et al. (2009) and energy dissipation Hack and Schmidt (2021) are induced by the same mechanism.
It is noteworthy that vortex-induced separation is associated with blow-up of solutions of the Prandtl boundary-layer equations, leading to diverging wall-normal velocities Peridier et al. (1991). This type of singularity was first identified numerically by Dommelen & Shen in the Prandtl problem of an impulsively accelerated cylinder van Dommelen and Shen (1980); Van Dommelen and Shen (1982) and the blow-up has since been rigorously established Kukavica et al. (2017); Collot et al. (2021). The boundary-layer equations cannot explain quantitatively all of the features of such extreme events, but blow-up of the Prandtl solutions is probably a necessary antecedent.
VII.2 Smooth Vortices Impinging on a Wall
Given the above observations, a promising situation in which to observe anomalous dissipation and breakdown of weak-strong uniqueness is the problem of compact vortices in an otherwise quiescent flow impinging on a solid wall. The numerical study of the simplest such example, a dipole pair of vortices impacting on a flat wall in two space dimensions, was pioneered by Orlandi Orlandi (1990) and studied subsequently at higher resolution Clercx and van Heijst (2002); Nguyen van yen et al. (2018). This example shares with the d’Alembert potential flow the very important feature that a smooth Euler solution exists globally in time, in this case because of the restriction to two space dimensions. The smooth Euler solution corresponds here to the pair of vortices hitting the wall and then, under the influence of their image vortices, propagating along the wall in opposite directions.
The latest numerical solutions Nguyen van yen et al. (2018) of the Navier-Stokes equation with this same initial data at very high Reynolds numbers show quite different behavior, with the vortices rebounding from the wall and inducing thereby a cascade of separation of vorticities of alternating signs. This high- solution is vividly compared with the smooth Euler solution in a supplementary movie of Nguyen van yen et al. (2018) at https://doi.org/10.1017/jfm.2018.396. The highest Reynolds number achieved by Nguyen van yen et al. (2018) was where is the initial maximum velocity and is a measure of the radius of the initial vortices. A special adaptive grid was used for the computations to assure that length scales of order could be accurately resolved in the vicinity of the wall. This simulation provides rather convincing prima facie evidence for co-existence of a smooth Euler solution and a quite distinct inviscid limit solution with the same initial data.
Most importantly, the paper Nguyen van yen et al. (2018) presented strong evidence also for the extreme events that are required in order to violate weak-strong uniqueness. Figure 12a of Nguyen van yen et al. (2018) plotted versus Reynolds number the maximum vorticity at the wall, which was found to scale as before the blow-up of the Prandtl solution but as a short time after the blow up. Since skin friction is related to wall vorticity by for stick b.c., the above observation corresponds to as at least pointwise. Furthermore, Figure 12b of Nguyen van yen et al. (2018) plotted also versus Reynolds number the enstrophy which was likewise found to scale as before the Prandtl blow-up but as afterward. Because total viscous energy dissipation can be written as the latter scaling implies a dissipative anomaly. No evidence was provided in Nguyen van yen et al. (2018) for violation of the Drivas-Nguyen conditions (IV.4), (IV.5), but the authors did note “… a blow-up of the wall-normal velocity associated with an abrupt acceleration of fluid particles away from the wall” (p.697). Although further confirmation would be desirable, the paper Nguyen van yen et al. (2018) presents to our knowledge the most complete evidence for physical violation of weak-strong uniqueness within the class of inviscid limits.
VII.3 Breakdown of Deterministic Navier-Stokes?
An important feature of the conditions required to violate weak-strong uniqueness is that they are so extreme that they threaten the validity of a macroscopic hydrodynamic description. To explain this, we may argue phenomenologically. Associated to local, instantaneous skin friction on a body surface, one can introduce a fluctuating viscous length
This length is analogous to the fluctuating dissipation scale considered in bulk turbulence Paladin and Vulpiani (1987); Schumacher (2007). Note that a strong momentum anomaly corresponds to which yields the “Kato length” However, using again the estimate from kinetic theory that then gives also
where is the Mach number. This is only larger than by and thus dangerously close to length-scales at which no hydrodynamic description can be accurate. Furthermore, thermal fluctuations become sizable already at lengths de Zarate and Sengers (2006); Bandak et al. (2022); Bell et al. (2022). Thus, a quantitatively correct description of such extreme wall events is probably not provided by deterministic Navier-Stokes equations but instead by some version of fluctuating hydrodynamics including fluid-solid friction effects Reichelsdorfer (2016).
What is crucial to emphasize is that these tiny lengths may arise not only for the case of impulsive acceleration, which is obviously very singular, but even for apparently much more regular flows, for example, with gradually accelerated smooth bodies or smooth dipolar vortices impinging on a flat wall. If the extreme events necessary to violate weak-strong uniqueness do not occur, then inviscid limits with smooth initial data necessarily coincide with the smooth Euler solution as long as that exists. A plausible mechanism to produce such extreme events is explosive separation of thin boundary-layers at extremely high Reynolds numbers.
VIII Conclusions
According to the mathematical results reviewed in this work, there are two main scenarios for very high Reynolds-number fluid flows interacting with solid walls, depending upon which of the following two limits holds: (i) or (ii)
(i) If then the infinite-Reynolds number limit coincides with the smooth Euler solution with the same initial data, as long as the latter exists. This is usually globally in time for potential flows. In that case, there is at most a weak energy dissipation anomaly with exact potential-flow initial data, although a strong dissipation anomaly might occur also if the initial potential flow is perturbed by turbulent, rotational fluctuations of small but non-vanishing energy.
(ii) If then the infinite-Reynolds number limit may be distinct from the smooth Euler solution with precisely the same initial data. A weak, singular Euler solution with a strong energy dissipation anomaly and vorticity cascade into the flow interior may coexist with the smooth Euler solution for the same initial data.
However, the condition requires very extreme near-wall events: non-zero energy dissipation in a thin “Kato layer” (viscous sub-layer & buffer layer) and discontinuity of the wall-normal velocity approaching the solid boundary through the inertial sublayer. The deterministic Navier-Stokes or any hydrodynamic description whatsoever may break down, at least locally within the spacetime vicinity of the extreme event.
Determining which of these possibilities is realized physically calls for a focused campaign of empirical investigation, both by numerical simulation and by laboratory experiment. Study of smoothly accelerated bodies looks especially promising, since this problem has obvious practical importance and yet exemplifies the fundamental issues. Computational efforts will need to take particular care for very fine space resolution near the body surface, for example with adaptive algorithms Chatzimanolakis et al. (2022), in order to capture (or rule out) the emergence of the requisite small-scales. Since the events of interest may occur only sporadically in space-time, computational techniques for sampling extreme events Farazmand and Sapsis (2017); Lestang et al. (2020); Rolland (2022) may be useful. Experiments will be challenging also because of the sporadic nature of the events of interest and the lack of resolution of conventional measurement tools, such as particle-imaging velocimetry, near the body surface. New methods of measurement currently being developed Örlü and Vinuesa (2020); Jassal and Schmidt (2023) might be crucial. Achieving high Reynolds numbers while maintaining hydraulic smoothness of the surface may require acceleration of large bodies. Laboratory experiments are, however, the only investigative tool that does not presuppose a particular mathematical model describing the flow and are thus indispensable.
These issues have importance beyond the specific problem of external flow past bluff bodies. Onsager’s analysis of the infinite Reynolds limit implies that observed anomalous dissipation of kinetic energy in high Reynolds turbulence requires suitable Hölder singularities of the limiting inviscid velocity field Onsager (1949); Eyink (2024a) and experiments have provided evidence that the predicted singularities indeed exist Lashermes et al. (2008); Debue et al. (2021); Faller et al. (2021). However, Onsager’s theory leaves unexplained the origin of these singularities. A common view has been that singularities arise from finite-time blow-up of a smooth Euler solution (Frisch (1995), §7.8). However, this mechanism is unavailable for 3D incompressible potential flows and for all 2D incompressible flows. Furthermore, such an origin of dissipative singularities would not account for the experimental observation that many flows exhibit anomalous dissipation only with hydraullically rough walls Nikuradse (1933); Cadot et al. (1997), in which pressure forces produce drag similarly as for bluff bodies. The conditional weak-strong uniqueness results for wall-bounded incompressible flows suggest a completely different origin of the required singularities: explosive events at the wall that lead in the inviscid limit to dissipative weak Euler solutions co-existing with the smooth Euler solution for the same initial-data. We remark again that the conditions that we established for weak-strong uniqueness in external flows past solid boidies are identical to those of Kelliher Kelliher (2017) for certain classes of internal flows. In fact, experimental studies in the specific example of the von Kármán flow show that many of the points of higest local energy cascade and vorticity magnitude appear in close vicinity to the rotating impellers that drive the flow (see Faller et al. (2021), Fig.5). It is reasonable to conjecture that the ultimate origin of dissipative singularities in high Reynolds-number incompressible fluid turbulence may lie in strong vorticity shed at solid walls during explosive separation.
Acknowledgements
We thank T. Drivas for stressing the significance of ref.Nguyen van yen et al. (2018). Our work was funded by the Simons Foundation, Targeted Grant MPS-663054 and Collaboration Grant MPS-1151713. G.E. thanks also the Department of Physics of the University of Rome ‘Tor Vergata’ for hospitality while this work was finalized and acknowledges support from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (Grant Agreement No. 882340).
Appendix A Navier-Slip Boundary Conditions
It is well-known that stick boundary conditions at the wall, on although widely adopted, are only an approximation to more accurate Navier slip b.c.
where is a friction coefficient (with units of velocity). In particular, such slip b.c. are necessary to model accurately a body impulsively accelerated through a molecular fluid, within a hydrodynamic framework Reddy (1967). However, the derivation of the Josephson-Anderson relation was previously explained using stick b.c. Eyink (2021, 2024b) and likewise its inviscid limit was demonstrated for those standard b.c. Quan and Eyink (2024). To apply our weak-strong uniqueness result with the slip b.c. (A.1), we must discuss briefly the slight changes required to those previous analyses.
First, we recall the standard result from kinetic theory Reichelsdorfer (2016); Jiang and Masmoudi (2017) obtained already by Maxwell Maxwell (1879), that slips arising from molecular effects have a slip length the mean-free-path length, and friction coefficient the sound speed. Because of the presence of the additional dimensional parameter (or ), usual Reynolds similarity breaks down. In fact, non-dimensionalizing the Navier-Stokes equations with large scale length and velocity (outer units) yields slip b.c.
where in addition to Reynolds number there appears also the Mach number which must be assumed sufficiently small for validity of the incompressible approximation. Since Reynolds similarity is broken, it now matters how the limit is achieved. In particular, increasing would eventually violate the condition so that we consider instead decreasing or especially increasing with fixed.
In mathematical parlance, we take dimensionless parameters and fixed, which is known as the case of “critical slip”. The theorems of Drivas and Nguyen (2018) still apply with Navier slip b.c., showing that inviscid limits yield dissipative weak Euler solutions under physically reasonable assumptions. Kato-type theorems have also been proved, at least for 2D domains Wang et al. (2017), implying strong convergence to the smooth Euler solution under conditions of vanishing dissipation in an -neighborhood of the boundary. Similar results have been proved without such conditions and assuming instead analytic initial data in a 2D flat wall geometry Tao et al. (2020); Nguyen (2020), but convergence holds only for a finite time over which analyticity is preserved. In fact, the case of a dipole vortex impinging on a 2D flat wall has been simulated numerically with critical Navier slip b.c. Nguyen van Yen et al. (2011) and anomalous dissipation seems to appear after sufficiently long times, with apparent breakdown of weak-strong uniqueness for inviscid limits.
The proof of the Josephson-Anderson relation Eyink (2021, 2024b), in fact, does not differ for stick and slip b.c., because the only boundary condition used in the derivation are The starting point is the force exerted on the body by the rotational flow,
where
is again a tangent vector field on the body surface. Standard arguments Lighthill (1986); Eyink (2021) yield the force-impulse relation and then Eyink (2024b) the JA-relation:
Integration by parts with yields
but it is no longer true that for slip b.c. In fact, a bit of computation shows that
where is the rank-2, symmetric matrix which defines the Weingarten map (or shape map) on the tangent space of the surface. Thus, for Navier slip b.c., there is an extra term in the JA-relation when written in this alternative form appropriate for taking the inviscid limit. Note however that the additional term vanishes as (or ) if uniformly in . This can be expected from simple energy estimates, since is fixed as
Finally, we note that the global balance for the rotational flow energy is easily computed from Eq.(3.17) in Eyink (2021) to be
where and (A.4) was used twice. The final term again is expected to vanish in the inviscid limit. The rigorous treaments of this limit in Quan and Eyink (2024, 2025a) carry through almost unchanged, yielding a version of (V.2) but with an additional dissipation term from surface friction. Thus, the condition again suffices to derive weak-strong uniqueness (or, if necessary, one can also include the assumption that the final term in (A.7) vanishes). However, since with in the limit it seems unlikely that